OpenStudy (anonymous):

What is the approximate volume of a cone with a height of 9 ft and radius of 3 ft? Use 3.14 to approximate pi, and express your final answer to the nearest tenth. @phi

OpenStudy (phi):

write down the formula for the volume of a cone then replace r with 3 and h with 9 can you do that ?

OpenStudy (anonymous):

V=1/3TTr^2h V=1/3TT(3)^2(9)

OpenStudy (phi):

now do the arithmetic.

OpenStudy (anonymous):

okay V=1/3TT*9*9 v= 1/3TT*81

OpenStudy (phi):

keep going

OpenStudy (anonymous):

okay um V=1/3*3.14*81 V=1/3*254.34 V=.33*254.34=83.9322

OpenStudy (anonymous):

@phi ?????

OpenStudy (anonymous):

would it be 84 if you rounded to the nerest tenth

OpenStudy (phi):

You have the right idea, but you have to be more careful. you changed 1/3 to 0.33 by .33 is not exactly 1/3 you should do 81/3 first what do you get?

OpenStudy (anonymous):

81/3 would equal ??? um would i have to divide 81 by 3

OpenStudy (phi):

your problem would be easier if you had done it this way: \[ \frac{1}{3} \cdot 3 \cdot 3 \cdot 9 \pi \] I wrote 3^2 as 3*3 next I would divide 3 evenly into 3 that leaves \[ 3 \cdot 9 \cdot \pi \] notice how dividing *first* gives you smaller numbers. now finish the problem

OpenStudy (phi):

btw, yes 81/3 means 81 divided by 3. you get 27

OpenStudy (anonymous):

okay so it would be 81ftTT^3

OpenStudy (phi):

? 81/3 pi or 27 pi

OpenStudy (anonymous):

okay

OpenStudy (phi):

now do 27*3.14

OpenStudy (anonymous):

so 84.78

OpenStudy (anonymous):

would then equal 85 right

OpenStudy (phi):

to round to the nearest 10th you would look at the 100th place. if 5 or bigger round up. here the 10th's place is 7 and the digit to the right is 8. that is 5 or bigger so round up 84.8

OpenStudy (anonymous):

okay